OpenStudy (anonymous):

could someone help me find the limit of (h^2 + 8h)/h as h approaches 0?

OpenStudy (anonymous):

I am not sure if it will be infinity or 8

OpenStudy (anonymous):

look at it again and let me ask you....what is 1/10000000000 approximately equal to? just say.....a really BIG number, or a really SMALL number

OpenStudy (anonymous):

a very small number.

OpenStudy (anonymous):

infinity

OpenStudy (anonymous):

sorry about that....first thing you should ALWAYS do is....plug the number in for h....what do you get when you plug 0 in for every "h" here?

OpenStudy (anonymous):

0/0

OpenStudy (anonymous):

right....is that a real answer? or is it undefined?

OpenStudy (anonymous):

undefined.

OpenStudy (anonymous):

correct....your answer would be...there is no limit as h approaches 0

OpenStudy (anonymous):

\[\frac{h^2+8h}{h}=\frac{h+8}{1}=h+8\]

OpenStudy (anonymous):

forgive me....@SithsAndGiggles is correct

OpenStudy (anonymous):

8